OpenStudy (anonymous):

To the nearest tenth of a percent, what is 8/13 written as a percent? A. 0.6% B. 6.2% C. 61.5% D. 615.4%

OpenStudy (yuliacosta111):

= 8/13 = 8 ÷ 13 = 0.615385 0.615385 x 100 = 61.5385%

OpenStudy (yuliacosta111):

its C